A body is projected along a rough horizontal road with speed 10m/s . If the coefficent of kinetic friction is 0.1 , how far will it  will travel before coming to rest ? 

Grade:9

1 Answers

Amar Kumar
10 years ago

just see

acceleration, a  =  - μg = - 1 m/s2

v2 = u2 + 2as ,  s = 100/2 = 50 m


Thanks & Regards 

Amar Kumar, 

askIITians Faculty

M.Tech, IIT Delhi